One solution contains 40% alcohol and another contains 72% alcohol. How much of each should be mixed together to obtain 16 ounces of a 62% alcohol solution?
Answer by fractalier(6550)

Let x be the amount of 40% alcohol. The 16-x must be the amount of 72% alcohol. The equation becomes, once you make percents into whole numbers...
x(40) + (16-x)(72) = 16(62)
40x + 1152 - 72x = 992
Simplifying gives
-32x = -160
and
x = 5
So 5 oz of 40% and 11 oz of 72% will do the trick.